Sulfuric acid can be produced by the following sequence of reactions.              S8 + 8 O2 --> 8 SO2           ...

  1. Sulfuric acid can be produced by the following sequence of reactions.

             S8 + 8 O2 --> 8 SO2

              2 SO2 + O2   --> 2 SO3

              SO3 + H2O --> H2SO4

You find a source of sulfur that is 65.0% sulfur by mass. How many grams of this starting material would be required to produce 87.0 g of sulfuric acid? Explain and show work.
Answer #1

Moles of sulfuric acid required = 87.0g/98.0785g/mol = 0.887045moles

Moles of SO3 required = moles of sulfuric acid = 0.887045

Moles of SO2 required = (2/2)×moles of SO3 = 0.887045

Moles of S8 = (1/8)×moles of SO2 = (1/8)×0.887045

= 0.11088mol

Mass of sulfur required = number of moles × molar mass = 0.11088mol × 256.52g/mol = 28.4431 g

Amount of starting material required = (100/65)×28.4431g

= 43.7586 g

= 43.8 g. (Answer)